OpenStudy (anonymous):

Help Please. Solve the equation below by completing the square and then solving for x. x^2+x+1=7/4

OpenStudy (anonymous):

x= 1/2 & x= -3/2

OpenStudy (anonymous):

Thanks @music101

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

i was gonna show u the work, but i guess u don't need it